can you please help me find the domain, range, intercepts and asymptotes for the function, f(x)=(3/(x+2))
Answer by stanbon(57361)

f(x)=(3/(x+2))
find the domain:
all Real numbers except x=-2
---------------
range: All Real numbers
------------------
intercepts:
x-intercepts
Let y=0 and solve for "x"
0=3/(x_2)
There is no solution, so there are no x-intercepts.
-----------
y-intercept:
Let x=0, then y=3/2
y-intercept at (0,3/2)
----------------
asymptotes
Vertical asymptote when the denominator is zero; at x=-2
Horizontal asymptote at y=0/1 = 0 (0 and 1 are the coefficients of x)
-----------
Cheers,
Stan H.